OpenStudy (anonymous):

A = Pert, where P is the initial investment and r is the interest rate. What rate of interest compounded continuously is needed for an investment of $500 to grow to $900 in 10 years? show work plz

OpenStudy (dumbcow):

well P = 500 A = 900 t =10 solve for r \[900 = 500e^{10r}\] divide by 500 \[e^{10r} = 9/5\] take natural log of both sides \[10r = \ln (9/5)\] divide by 10 \[r = \frac{\ln (9/5)}{10}\]

OpenStudy (dumbcow):

did you follow that?

OpenStudy (anonymous):

Theres suppose to be a simplier answer apparently and I dont get how you take in9/5 --- 10

OpenStudy (dumbcow):

just put it in your calculator ln(9/5) = .5877

OpenStudy (dumbcow):

im guessing they want it as a percent multiply r by 100 then

OpenStudy (anonymous):

thank you so much!